Shift report 20000704 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Jul 04 2000 - 00:01:08 EDT

  • Next message: http account: "Shift report 20000704 00:00-08:00"

    ------------------------------------------------------------
    Date: Tue Jul  4 00:01:07 2000, Shift: 16:00-24:00
    Supervisor: ko, with: bs, tv
    Log book#2 pages 126 - 127
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine was up since 18:00, two good stores- at 18:00
       two hours, at 21:30 still going strong with
       blue=680 and yellow=314.
    
    2. Two good data runs. Intensity and trigger rates look
       higher than yesterday's stores. Summary plots have
       changed- plots removed and added.
    
    3. T2 is sick, T2B anode tripped once.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Jul 04 2000 - 00:05:15 EDT